### v2

# Stable Flow: Vital Layers for Training-Free Image Editing

Abstract
Diffusion models have revolutionized the field of content synthesis and editing. Recent models have replaced the traditional UNet architecture with the Diffusion Transformer (DiT), and employed flow-matching for improved training and sampling. However, they exhibit limited generation diversity. In this work, we leverage this limitation to perform consistent image edits via selective injection of attention features. The main challenge is that, unlike the UNet-based models, DiT lacks a coarse-to-fine synthesis structure, making it unclear in which layers to perform the injection. Therefore, we propose an automatic method to identify “vital layers” within DiT, crucial for image formation, and demonstrate how these layers facilitate a range of controlled stable edits, from non-rigid modifications to object addition, using the same mechanism. Next, to enable real-image editing, we introduce an improved image inversion method for flow models. Finally, we evaluate our approach through qualitative and quantitative comparisons, along with a user study, and demonstrate its effectiveness across multiple applications.
Figure 1: Stable Flow. Our training-free editing method is able to perform various types of image editing operations, including non-rigid editing, object addition, object removal, and global scene editing. These different edits are done using the same mechanism.
Project page is available at: https://omriavrahami.com/stable-flow This research was performed while Omri was at Snap.
1 Introduction
Over the recent years, we have witnessed an unprecedented explosion in creative applications of generative models, fueled by diffusion-based models [79, 81, 37, 80]. Recent models, such as FLUX [47] and SD3 [26], have replaced the traditional UNet architecture [74] with the Diffusion Transformer (DiT) [64], and adopted flow matching [51, 50, 6] as a superior alternative for training and sampling.
These flow-based models are based on optimal transport conditional probability paths, resulting in faster training and sampling, compared to diffusion models. This is attributed [50] to the fact that they follow straight line trajectories, rather than curved paths. One of the known consequences of this difference, however, is that these models exhibit lower diversity than previous diffusion models [27], as shown in Figure 2(1-2). While reduced diversity is generally considered an undesirable characteristic, in this paper, we suggest leveraging it for the task of training-free image editing, as shown in Figure 2(3) and Figure 1.
Specifically, we explore image editing via parallel generation [91, 19, 5], where features from the generative trajectory of the source (reference) image are injected into the trajectory of the edited image. Such an approach has been shown effective in the context of convolutional UNet-based diffusion models [19], where the roles of the different attention layers are well understood. However, such understanding has not yet emerged for DiT [64]. Specifically, DiT does not exhibit the same fine-coarse-fine structure of the UNet [64], hence it is not clear which layers should be tampered with to achieve the desired editing behavior.
To address this gap, we analyze the importance of the different components in the DiT architecture, in order to determine the subset that should be injected while editing. More specifically, we introduce an automatic method for detecting a set of vital layers — layers that are essential for the image formation — by measuring the deviation in image content resulting from bypassing each layer. We show that there is no simple relationship between the vitality of a layer and its position in the architecture, i.e., the vital layers are spread across the transformer.
A close examination of the vital layers suggests that the injection of features into these layers strikes a good balance in the multimodal attention between the reference image content and the editing prompt. Consequently, limiting the injection of features to only the vital layers tends to yield a stable edit, i.e., an edit that changes only the part(s) specified by the text prompt, while leaving the rest of the image intact, as demonstrated in Figure 2(3). We demonstrate that performing the same feature injection, enables performing a variety of image edits, including non-rigid editing, addition of objects, and scene changes, as demonstrated in Figure 1.
(1) SDXL [67]
Refer to caption Refer to caption Refer to caption Refer to caption
(2) FLUX [47]
Refer to caption Refer to caption Refer to caption Refer to caption
(3) Stable Flow
Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a + “dog wearing + “cat wearing + “casting
dog and a cat…” a blue hat” yellow glasses” shadows”
Figure 2: Leveraging Reduced Diversity. Using the same initial seed with different editing prompts, diffusion models such as (1) SDXL generate diverse results (different identities of the dog and the cat), while (2) FLUX generates a more stable (less diverse) set of results out-of-the-box. However, there are still some unintended differences (the dog is standing in the leftmost column and sitting in the others, the color of the cat is changing, and the road is different on the right). Using our approach, (3) Stable Flow, the edits are stable, maintaining consistency of the unrelated content.
In order to support editing real images, it is typically necessary to invert them first [80, 54]. We employ an Inverse Euler Ordinary Differential Equation (ODE) solver to invert real images in the FLUX model [47]. However, this method fails to reconstruct the input image in a satisfactory manner. To improve reconstruction accuracy, we introduce an out-of-distribution (OOD) nudging technique, in which we apply a small scalar perturbation to the clean latent before inverting it. We demonstrate that this makes FLUX less prone to undesired changes in the image during the forward pass.
Finally, we compare our method against its baselines qualitatively and quantitatively, and reaffirm the results with a user study. We also demonstrate several applications of our method.
In summary, our contributions are: (1) we propose an automatic method to detect the set of vital layers in DiT models and demonstrate how to use them for image editing; (2) we are the first method to harness the limited diversity of flow-based models to perform different image editing tasks using the same mechanism; and (3) we present an extension that allows editing real images using the FLUX model.
Figure 3: Layer Removal. (Left) Text-to-image DiT models consist of consecutive layers connected through residual connections [34]. Each layer implements a multimodal diffusion transformer block [26] that processes a combined sequence of text and image embeddings. (Right) For each DiT layer, we perform an ablation by bypassing the layer using its residual connection. Then, we compare the generated result on the ablated model with the complete model using a perceptual similarity metric.
2 Related Work
Text-Driven Image Editing.
After the emergence of text-to-image models, many works have suggested using them for various applications [8, 38, 21, 13, 76], including text-driven image editing tasks [40, 66]. SDEdit [53] addressed the image-to-image translation task adding noise and denoising with a new prompt. Blended Diffusion [9, 10] suggested performing localized image editing [39, 48, 16, 56, 63] incorporating an input mask into the diffusion process in a training-free manner, while GLIDE [55], ImagenEditor [89], and SmartBrush [93] offered fine-tuning the model on a given input mask. Other works suggested inferring the mask from the input image [22, 88] or via an additional click input [70]. Other works [35, 84, 61, 19] offered to inject information from the input image using parallel generation. While some methods [14, 46, 85, 100, 99] suggested fine-tuning the model per-image, a recent line of work suggested training a designated model on a large synthetic dataset for instruction-based edits [18, 97, 77]. However, none of the above methods is a training-free method that supports non-rigid editing, object adding/replacement and scene editing altogether.
Image Inversion.
In the realm of generative models, inversion [92] is the task of finding a code within the latent space of a generator [32, 44, 45] that faithfully reconstructs a given image. Initial methods were developed for GAN models [1, 2, 102, 71, 83, 65, 101, 4, 72, 15, 25, 60, 95], and more recently for diffusion-based models [80, 54, 87, 59, 41, 24, 30, 33, 52, 17]. In this work, we suggest inverting a real image in flow models using latent nudging, as we found that the standard inverse ODE solver is insufficient.
3 Method
Our goal is to edit images based on text prompts while faithfully preserving the unedited regions of the source image. Given an input image $x$ and an editing prompt $p$, we aim to generate a modified image $\hat{x}$ that exhibits the desired changes specified by $p$ while maintaining the original content elsewhere. We leverage the limited diversity of the FLUX model and further constrain it to enable such stable image edits through selective injection of attention features of the source image into the process that generates $\hat{x}$. Our approach is described in more detail below. In Section 3.1, we evaluate layer importance in the DiT model by analyzing the perceptual impact of layer removal (Figure 3). Next, in Section 3.2, we employ the most influential layers (termed vital layers) for image editing through attention injection. Finally, in Section 3.3, we extend our method to real image editing by inverting images into the latent space using the Euler inverse ODE solver, enhanced by latent nudging.
3.1 Measuring the Importance of DiT Layers
Recent text-to-image diffusion models [37, 69, 73, 75, 67] predominantly use CNN-based UNets, which exhibit well-understood layer roles. In discriminative tasks, early layers detect simple features like edges, while deeper layers capture higher-level semantic concepts [96, 78]. Similarly, in generative models, early-middle layers determine shape and color, while deeper layers control finer details [44]. This structure has been successfully exploited in text-driven editing [84, 19, 31] through targeted manipulation of UNet decoder layers [74].
In contrast, state-of-the-art text-to-image DiT [64] models (FLUX [47] and SD3 [26]) employ a fundamentally different architecture, as shown in Figure 3(left). These models consist of consecutive layers connected through residual connections [34], without convolutions. Each layer implements a multimodal diffusion transformer block [26] (MM-DiT-Block) that processes a combined sequence of text and image embeddings. Unlike in UNets, the roles of the different layers are not yet intuitively clear, making it challenging to determine which layers are best suited for image editing.
To quantify layer importance in the FLUX model, we devised a systematic evaluation approach. Using ChatGPT [57], we automatically generated a set $P$ of $k=64$ diverse text prompts, and draw a set $S$ of random seeds. Each of these prompts was used to generate a reference image, yielding in a set $G_{\textit{ref}}$. For each DiT layer $\ell\in\mathbb{L}$, we performed an ablation by bypassing the layer using its residual connection, as illustrated in Figure 3(right). This process generated a set of images $G_{\ell}$ from the same prompts and seeds. See the supplementary material for more details.
To assess the impact of each layer, we measured the perceptual similarity between $G_{\textit{ref}}$ and $G_{\ell}$ and using DINOv2 [58] (see Figure 3). The results, plotted in Figure 4, show that removing certain layers significantly affects the generated images, while others have minimal impact. Importantly, influential layers are distributed across the transformer rather than concentrated in specific regions. We formally define the vitality of layer $\ell$ as:
$vitality(\ell)=1-\frac{1}{k}\sum_{s\in S,p\in P}d(M_{\text{full}}(s,p),M_{\text{-}\ell}(s,p)),$ (1)
where $M_{\text{full}}$ represents the complete model, $-M_{\text{-}\ell}$ denotes the model with layer $\ell$ omitted, and $d(\cdot,\cdot)$ is the perceptual similarity metric. The set of vital layers $V$ is then defined as:
$V=\left\{\ell\in\mathbb{L}\mid vitality(\ell)\geq\tau_{vit}\right\},$ (2)
where $\tau_{vit}$ is the vitality threshold.
Figure 5 illustrates the qualitative differences between vital and non-vital layers. While bypassing non-vital layers results in minor alterations, removing vital layers leads to significant changes: complete noise generation ($G_{0}$), global structure and identity changes ($G_{18}$), and alterations in texture and fine details ($G_{56}$).
00555510101010151515152020202025252525303030303535353540404040454545455050505055555555606060600.80.80.80.80.850.850.850.850.90.90.90.90.950.950.950.95Layer IndexPerceptual Similarity
Figure 4: Layer Removal Quantitative Comparison. As explained in Section 3.1, we measured the effect of removing each layer of the model by calculating the perceptual similarity between the generated images with and without this layer. Lower perceptual similarity indicates significant changes in the generated images (Figure 5). As can be seen, removing certain layers significantly affects the generated images, while others have minimal impact. Importantly, influential layers are distributed across the transformer rather than concentrated in specific regions. Note that the first vital layers were omitted for clarity (as their perceptual similarity approached zero).
$G_{\textit{ref}}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{0}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{5}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{18}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{52}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{56}$
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Layer Removal Qualitative Comparison. As explained in Section 3.1, we illustrate the qualitative differences between vital and non-vital layers. While bypassing non-vital layers ($G_{5}$ and $G_{52}$) results in minor alterations, bypassing vital layers leads to significant changes: complete noise generation ($G_{0}$), global structure and identity changes ($G_{18}$), and alterations in texture and fine details ($G_{56}$).
Figure 6: Multi-Modal Attention Distribution. Given an input image of a man, we edit it to hold an avocado by injecting the reference image attention activations in the vital layers only (left), or in the non-vital layers (right), and visualize the multimodal attention of two points: a yellow point in a region that should remain unchanged (requiring copying from the reference image), and a red point in an area targeted for editing (requiring generation based on the text prompt). As can be seen, in vital layers (left), points meant to remain unchanged show dominant attention to visual features, while points targeted for editing exhibit stronger attention to relevant text tokens (e.g., “avocado”). Conversely, non-vital layers (right) show predominantly image-based attention even in regions marked for editing. This suggests that injecting features into vital layers strikes a good multimodal attention balance between preserving source content and incorporating text-guided modifications.
3.2 Image Editing using Vital Layers
Given a source image $x$ generated with a known seed $s$ and prompt $p$, we aim to modify $p$ and generate an edited image $\hat{x}$ that exhibits the desired changes, while otherwise preserving the source content. We adapt the self-attention injection mechanism, previously shown effective for image and video editing [91, 19] in UNet-based diffusion models, to the DiT-based FLUX architecture. Since each DiT layer processes a sequence of image and text embeddings, we propose generating both $x$ and $\hat{x}$ in parallel while selectively replacing the image embeddings of $\hat{x}$ with those of $x$, but only within the vital layers set $V$.
Remarkably, as shown in Figure 1, this training-free approach successfully performs diverse editing tasks, including non-rigid deformations, object addition, object replacement, and global modifications, all using the same set of vital layers $V$.
To understand this effectiveness, we analyze the multimodal attention patterns in FLUX. Each visual token simultaneously attends to all visual and text tokens, with attention weights normalized across both modalities. Figure 6 contrasts the attention patterns in vital versus non-vital layers at two key points: a yellow point in a region that should remain unchanged (requiring copying from the reference image), and a red point in an area targeted for editing (requiring generation based on the text prompt). In vital layers (left), points meant to remain unchanged show dominant attention to visual features, while points targeted for editing exhibit stronger attention to relevant text tokens (e.g., “avocado”). Conversely, non-vital layers (right) show predominantly image-based attention even in regions marked for editing. This suggests that injecting features into vital layers strikes a good multimodal attention balance between preserving source content and incorporating text edits.
3.3 Latent Nudging for Real Image Editing
(a) w/o nudging
Refer to caption Refer to caption Refer to caption
(b) w nudging
Refer to caption Refer to caption Refer to caption
Input image Reconstruction “Raising its hand”
Figure 7: Latent Nudging. As described in Section 3.3, when inverting a real image, (a) a simple inverse Euler ODE solver leads to corrupted image reconstructions and unintended modifications during editing. On the other hand, (b) using our latent nudging technique significantly reduces reconstruction errors and better constrains edits to the intended regions.
[…]
To edit real images, we must first invert them into the latent space, transforming samples from $p_{1}$ to $p_{0}$. We initially implemented an inverse Euler ODE solver for FLUX by reversing the vector field prediction. Given the forward Euler step:
$z_{t-1}=z_{t}+(\sigma_{t+1}-\sigma_{t})*u_{t}(z_{t})$ (3)
where $z_{t}$ represents the latent at timestep $t$, $\sigma_{t}$ is the optimal transport standard deviation at time $t$, and $u_{t}$ is the learned vector field, we proposed the inverse step:
$z_{t}=z_{t-1}+(\sigma_{t}-\sigma_{t+1})*u_{t}(z_{t-1})$ (4)
assuming $u_{t}(z_{t})\approx u_{t}(z_{t-1})$ for small steps.
However, as Figure 7(a) demonstrates, this approach proves insufficient for FLUX, resulting in corrupted image reconstructions and unintended modifications during editing. We hypothesize that the assumption $u(z_{t})\approx u(z_{t-1})$ does not hold, which causes the model to significantly alter the image during the forward process. To address this, we introduce latent nudging: multiplying the initial latent $z_{0}$ by a small scalar $\lambda=1.15$ to slightly offset it from the training distribution. While this modification is visually imperceptible (Figure 7(b)), it significantly reduces reconstruction errors and constrains edits to the intended regions. See the supplementary material for more details and ablations.
4 Experiments
In Section 4.1 we compare our method against its baselines, both qualitatively and quantitatively. Next, in Section 4.2 we conduct a user study and report results. Furthermore, in Section 4.3 we present the ablation study results. Finally, in Section 4.4 we demonstrate several applications.
4.1 Qualitative and Quantitative Comparison
Input SDEdit [53] P2P+NTI [35, 54] Instruct-P2P [18] MagicBrush [97] MasaCTRL [19] Stable Flow (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“The cat is yelling and raising its paw”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A rabbit toy sitting and wearing pink socks during the late afternoon
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A rubber duck next to a purple ball during a sunny day”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A dog with a small collar lifting its paw while wearing red glasses
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A bottle next to an apple. There is a heart painting on the wall.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A doll with a green body wearing a hat
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A man with a long hair
Figure 8: Qualitative Comparison. We compare our method on real images against the baselines. SDEdit [53] faces challenges with preserving object identities and backgrounds (e.g., rabbit and cat examples). P2P+NTI  [35, 54] struggles with both preserving object identities (e.g., rabbit and lion dolls examples) and adding new objects (e.g., missing ball in the duck example and missing heart in the bottle example). Instruct-P2P [18] and MagicBrush [97] struggle with performing non-rigid editing (e.g., raising of the paws in dog and cat examples, and the sitting of the rabbit in its example). MasaCTRL [19] has difficulty with preserving object identities (e.g., cat, dog and lion doll examples) and adding new objects (e.g., missing ball in the duck example and missing socks in the rabbit example). Our method, on the other hand, is able to adhere to the editing prompt while preserving the identities.
Table 1: Quantitative Comparison. We compare our method against the baselines in terms of text similarity ($\text{CLIP}_{txt}$), image similarity ($\text{CLIP}_{img}$) and image-text direction similarity ($\text{CLIP}_{dir}$). As can be seen, P2P+NTI [35, 54], Instruct-P2P [18], and MasaCTRL [19] suffer from low similarity to the text prompt. SDEdit [94] and MagicBrush [97] adhere more to the text prompt, but they struggle with image similarity and image-text direction similarity. Our method, on the other hand, achieves better image and image-text direction similarity.
Method $\text{CLIP}_{txt}$ ()($\uparrow$)( ↑ ) $\text{CLIP}_{img}$ ()($\uparrow$)( ↑ ) $\text{CLIP}_{dir}$ ()($\uparrow$)( ↑ ) SDEdit [94] 0.24 0.71 0.07 P2P+NTI [35, 54] 0.21 0.76 0.08 Instruct-P2P [18] 0.22 0.87 0.07 MagicBrush [97] 0.24 0.88 0.11 MasaCTRL [19] 0.20 0.76 0.03 Stable Flow (ours) 0.23 0.92 0.14
Table 2: Ablation Study. We conduct an ablation study and find that performing attention injection in all the layers or performing an attention extension in all the layers significantly reduces the text similarity. Furthermore, performing an attention injection in the non-vital layers or removing the latent nudging reduces the image similarity.
Method $\text{CLIP}_{txt}$ ()($\uparrow$)( ↑ ) $\text{CLIP}_{img}$ ()($\uparrow$)( ↑ ) $\text{CLIP}_{dir}$ ()($\uparrow$)( ↑ ) Stable Flow (ours) 0.23 0.92 0.14 Injection all layers 0.17 0.98 0.00 Injection non-vital layers 0.25 0.72 0.09 Extension all layers 0.18 0.98 0.01 w/o latent nudging 0.22 0.62 0.05
Table 3: User Study. We compare our method against the baselines using the standard two-alternative forced-choice format. Users were asked to rate which editing result is better (Ours vs. the baseline) in terms of: (1) target prompt adherence, (2) input image preservation, (3) realism and (4) overall edit quality. We report the win rate of our method compared to each baseline. As shown, our method outperforms the baselines across all categories, achieving a win rate higher than the random chance of 50%.
Ours vs Prompt Adher. ()($\uparrow$)( ↑ ) Image Pres. ()($\uparrow$)( ↑ ) Realism ()($\uparrow$)( ↑ ) Overall ()($\uparrow$)( ↑ ) SDEdit [53] 69.00% 68.00% 63.66% 70.66% P2P+NTI [35, 54] 76.00% 71.00% 72.66% 65.33% Instruct-P2P [18] 76.33% 75.66% 68.00% 60.33% MagicBrush [97] 61.33% 67.33% 76.66% 74.00% MasaCTRL [19] 82.33% 80.00% 80.33% 72.00%
We compare our method against the most relevant text-driven image editing methods. We re-implement SDEdit [53] using the FLUX.1-dev [47] model, and use the official public implementations of P2P+NTI [35, 54], Instruct-P2P [18], MagicBrush [97] and MasaCTRL [19]. See the supplementary material for more details.
In Figure 8 we compare our method against the baselines qualitatively on real images. As can be seen, SDEdit [53] has difficulty maintaining object identities and backgrounds. P2P+NTI  [35, 54] struggles with preserving object identities and with adding new objects. Instruct-P2P [18] and MagicBrush [97] face challenges with non-rigid editing. MasaCTRL [19] struggles with preserving object identities and adding new objects. Our method, on the other hand, is adhering to the editing prompt while preserving the identities.
To quantify the performance of our method and the baselines, we prepared an evaluation dataset based on COCO [49], that in contrast to previous benchmarks [77, 97], also contains non-rigid editing tasks. We start by filtering the dataset automatically to contain at least one prominent non-rigid body. Next, for each image, we use various image editing tasks (non-rigid editing, object addition, object replacement and scene editing) that take into account the prominent object, resulting in a total dataset of 3,200 samples. See the supplementary material for more details.
We evaluated the editing results using three metrics: (1) $\text{CLIP}_{img}$ that measures the similarity between the input image and the edited image, (2) $\text{CLIP}_{txt}$ that measure the similarity between the edited image and the target editing prompt, and (3) $\text{CLIP}_{dir}$ [62, 29] that measures the similarity between the direction of the prompt change and the direction of the image change.
As can be seen in Table 1, P2P+NTI [35, 54], Instruct-P2P [18], and MasaCTRL [19] suffer from low similarity to the text prompt. SDEdit [94] and MagicBrush [97] adhere more to the text prompt, but they struggle with image similarity and image-text direction similarity. Our method, on the other hand, is able to achieve better image and image-text direction similarity.
4.2 User Study
We conduct an extensive user study using the Amazon Mechanical Turk (AMT) [7] platform, with the automatically generated test examples from Section 4.1. We compare all the baselines against our method using standard two-alternative forced-choice format. Users were given the input image, the edit text prompt, and two editing results (one from our method and one from the baseline). For each comparison, the users were asked to rate which editing result is better in terms of: (1) target prompt adherence, (2) input image preservation, (3) realism and (4) overall edit quality (i.e., when taking all factors into account). As can be seen in Table 3, our method is preferred over the baselines in overall terms as well as the other terms. See the supplementary material for more details and statistical analysis.
4.3 Ablation Study
We conduct an ablation study for the following cases: (1) Attention injection in all layers — we perform the attention injection that is described in Section 3.2 in all the layers (instead on the vital layers only). (2) Attention injection non-vital layers — we performed the attention injection in some non-vital layers (same amount of layers as vital layers). (3) Attention extension — instead of performing attention injection as described in Section 3.2, we extended [36, 82] the attention s.t. the generated images can attend to the reference image, as well as themselves. (4) w/o latent nudging — we omitted the latent nudging component (Section 3.3).
As can be seen in Table 2, we found that (1) performing attention injection in all the layers or performing (3) an attention extension in all the layers, significantly harms the text similarity. In addition, (2) performing an attention extension in the non-vital layers or (4) removing the latent nudging reduces the image similarity.
4.4 Applications
(1) Inc. Editing
Refer to caption Refer to caption Refer to caption Refer to caption
Input “Holding hands” “Wearing glasses” “Next to an albino
porcupine”
(2) Const. Style
Refer to caption Refer to caption Refer to caption Refer to caption
Input “Statue of Liberty” “Taj Mahal” “Eiffel Tower”
(3) Text Editing
Refer to caption Refer to caption Refer to caption Refer to caption
Input “Man holding a “Man holding a “Man holding a
sign with the sign with the sign with the
text ‘diffusion’ uppercase text text ‘Flow’ ”
in a blue color” ‘DIFFUSION’ ”
Figure 9: Applications. Our method can be used for various applications: (1) Incremental Editing — starting from a scene of two kids, the user can refine the image iteratively by making the kid hold hands, then wear glasses and finally add a porcupine next to them. (2) Consistent Style — starting from a scene with a given style, such as an animation of the Great Pyramid of Giza, the user can generate images of different places with the same style. (3) Text Editing — given a scene that contains text, our method is able to perform text-related editing such as color change, case change and text replacement.
As demonstrated in Figure 9, our method can be used for various applications: (1) Incremental Editing — starting from a given scene, the user can refine the image iteratively in a step-by-step manner. (2) Consistent Style — starting from a scene with a given style, the user can generate other images in the same style [28, 36, 42]. (3) Text Editing — given a scene that contains text, our method is able to perform text-related editing such as color change, case change and text replacement.
5 Limitations and Conclusions
(1) Style Editing
Refer to caption Refer to caption Refer to caption Refer to caption
Input “Animation” “Pencil sketch” “Oil painting”
(2) Object Drag.
Refer to caption Refer to caption Refer to caption Refer to caption
Input “On the right side “On the left side “On the bottom
of the frame” of the frame” of the frame”
(3) Bg. Replace
Refer to caption Refer to caption Refer to caption Refer to caption
Input “In the forest” “In the desert” “On the moon”
Figure 10: Limitations. Our method suffers from the following limitations: (1) Style Editing — given a photorealistic image of a boy, our method struggles with changing its style to an animation (the identity of the boy also changes), to pencil sketch (changes only to black&white) or to an oil painting (mainly makes the image smoother). (2) Object Dragging — given an image of a cat, our method is unable to drag it into different locations in the image, but changes the gaze of the cat instead. (3) Background Replacement — given an image of a rat on the road, our method unable to replace its background entirely (the road leaks).
As demonstrated in Figure 10, our method suffers from the following limitations: (1) Style Editing — given an input image in one style (e.g., photorealistic), our method struggles with changing it to a different style (e.g., oil painting), as it relies on attention injection (Section 3.2). (2) Object Dragging — given an image with an object, our method is unable to drag it [12] into different locations in the image, as text-to-image models often struggle [11] with spatial prompt adherence. (3) Background Replacement — given an input image, our method struggles with replacing its background entirely with no leakage [23].
In conclusion, we present Stable Flow, a training-free method for image editing that enables various image editing tasks using the attention injection of the same vital layers group. We believe that our fully-automated approach of detecting vital layers may be also beneficial for other use-cases, such as generative models pruning and distillation. We hope that our layer analysis will inspire more work in the field of generative models and expanding the possibilities for creative expression.
Acknowledgments. We thank Omer Dahary for his valuable help and feedback. This work was supported in part by the Israel Science Foundation (grants 1574/21 and 2203/24).
Supplementary Material
Appendix A Implementation Details
In Section A.1, we start by providing implementation details for our method. Next, in Section A.2, we provide the implementation details for the baselines we compared our method against. Later, in Section A.3, we provide the implementation details for the automatic evaluations dataset and metrics. Finally, in Section A.4 we provide the full details of the user study we conducted.
A.1 Method Implementation Details
Figure 11: Attention Injection. We adapt the self-attention injection mechanism, previously shown effective for image editing in UNet-based diffusion models, to the DiT-based FLUX architecture. Since each DiT layer processes a sequence of image and text embeddings, we propose generating both the reference image $x$ and generated image $\hat{x}$ in parallel while selectively replacing the attention keys and values that correspond to the image embeddings of $\hat{x}$ with those of $x$. This replacement is performed only within the vital layers set.
As described in Section 3.1, we started by collecting a dataset of $k=64$ text prompts using ChatGPT [57]. We instructed it to generate text prompts describing a diverse set of objects in different environments, with the focus on one main object. Then, we sampled $k$ seeds denoted by $S$ and used them to generate $k$ corresponding images $G_{\textit{ref}}$. Next, for each layer $l$, we bypass it by taking only the residual connection values. For each bypass, we generate $k$ images using the same seed set $S$ demoted by $G_{l}$. All the images were generated using Euler sampler in 15 steps and a guidance scale of $3.5$.
Next, to evaluate the effect of each layer $l$ on the final result, we compared the generated images $G_{l}$ with their corresponding images $G_{\textit{ref}}$ using the DINOv2 [58] perceptual similarity metric. We term the layers that effect the generated image the most (i.e., the layers with the lowest perceptual similarity) as vital layers, while the rest of the layers as non-vital layers. We found that the vital layers in the FLUX.1-dev model [47] are $[0,1,2,17,18,25,28,53,54,56]$. For visualization results, please refer to Section B.6. We empirically found that layer 2 can be removed from this set. In addition, the vital layers for the Stable Diffusion 3 (SD3) [26] model vital layers are: $[0,7,8,9]$. For more details, please refer to Section B.7.
In addition, as mentioned in Section 3.2, We adapt the self-attention injection mechanism, previously to be effective for image and video editing [91, 19] in UNet-based diffusion models, to the DiT-based FLUX architecture. Since each DiT layer processes a sequence of image and text embeddings, we propose generating both the reference image $x$ and generated image $\hat{x}$ in parallel while selectively replacing the image embeddings of $\hat{x}$ with those of $x$, but only within the vital layers set. A full visualization can be found in Figure 11.
Lastly, the variance list of the perceptual similarity of the different layers, as explained in Section 3.1, is as follows: [0.222, 0.041, 0.076, 0.08, 0.123, 0.101, 0.135, 0.124, 0.112, 0.105, 0.097, 0.12, 0.118, 0.086, 0.116, 0.067, 0.065, 0.116, 0.146, 0.065, 0.098, 0.061, 0.076, 0.077, 0.072, 0.086, 0.069, 0.067, 0.081, 0.091, 0.074, 0.062, 0.061, 0.044, 0.04, 0.054, 0.036, 0.038, 0.037, 0.04, 0.066, 0.04, 0.034, 0.044, 0.044, 0.031, 0.033, 0.036, 0.03, 0.032, 0.026, 0.026, 0.026, 0.079, 0.039, 0.037, 0.026].
A.2 Baselines Implementation Details
As explained in Section 4.1, we compare our method against the following baselines: SDEdit [53], P2P+NTI [35, 54], Instruct-P2P [18], MagicBrush [97], and MasaCTRL [19]. We reimplement SDEdit using the FLUX.1-dev model [47], and use the official implementation for the rest of the baselines.
We adapt the text prompts based on the baseline type: for SDEdit [53], P2P+NTI [35, 54], and MasaCTRL [19], we used the standard text prompt describing the desired edited scene (e.g., “A photo of a man with a red hat”). For the instruction-based baselines Instruct-P2P [18] and MagicBrush [97] we adapted the style to fit an instructional format (e.g., “Make the person wear a red hat”).
We used the following third-party implementations in this project:
A.3 Automatic Metrics Implementation Details
As explained in Section 4.1, we prepare an evaluation dataset based on the COCO [49] validation dataset. We begin by filtering the dataset automatically to include at least one prominent non-rigid body. More specifically, we filter only images containing humans or animals that at least one of them is prominent enough, but not too small, i.e., the prominent non-rigid body occupies at least 5% of the image but no more than 33%. Next, for each image, we apply various image editing tasks (non-rigid editing, object addition, object replacement, and scene editing) that take into account the prominent object from a list of different combinations, resulting in a total dataset of 3,200 samples. Examples of images from this dataset can be seen in Figure 14.
We evaluate the editing results using three metrics: (1) $\text{CLIP}_{img}$ which measures the similarity between the input image and the edited image by calculating the normalized cosine similarity of their CLIP image embeddings. (2) $\text{CLIP}_{txt}$ which measures the similarity between the edited image and the target editing prompt by calculating the normalized cosine similarity between the CLIP image embedding and the target text CLIP embedding. (3) $\text{CLIP}_{dir}$ [62, 29] which measures the similarity between the direction of the prompt change and the direction of the image change.
A.4 User Study Details
Figure 12: User Study Instructions. We provide the complete instructions for the user study we conducted using Amazon Mechanical Turk (AMT) [7] to compare our method with each baseline.
Figure 13: User Study Trial. We provide an example of a trial task in the user study conducted using Amazon Mechanical Turk (AMT) [7]. Users were asked four questions of a two-alternative forced-choice format. Complete instructions are shown in Figure 12.
Table 4: User Study Statistical Significance. A binomial statistical test of the user study results suggests that our results are statistically significant (p-value $<5\%$).
Ours vs Prompt Adher. Image Pres. Realism Overall p-value p-value p-value p-value SDEdit [53] $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-6}$ $<1\mathrm{e}{-8}$ P2P+NTI [35, 54] $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-8}$ $<6\mathrm{e}{-8}$ Instruct-P2P [18] $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-8}$ $<2\mathrm{e}{-4}$ MagicBrush [97] $<5\mathrm{e}{-5}$ $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-8}$ MasaCTRL [19] $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-8}$ $<1\mathrm{e}{-8}$
As described in Section 4.2 we conducted an extensive user study using the Amazon Mechanical Turk (AMT) [7] platform, using automatically generated test examples, as explained in Section A.3. We compared all the baselines with our method using a standard two-alternative forced-choice format. The users were given full instructions, as can be seen in Figure 12. Then, for each study trial, as shown in Figure 13, users were presented with an image and an instruction “Given the following input image of a {CATEGORY} where {CATEGORY} is the COCO category of the prominent object. The users were given two editing results — one from our method and one from the baseline, and were asked the following questions:
  1. 1.
    “Which of the results is better in adhering to the text prompt {PROMPT}?”, where {PROMPT} is the editing target prompt.
  2. 2.
    “Which of the results is better in preserving the information of the input image?”
  3. 3.
    “Which of the results looks more realistic?”
  4. 4.
    “Which of the results is better in overall?”
We collected five ratings per sample, resulting in 320 ratings per baseline, for a total of 1,920 responses. The time allotted per task was one hour, to allow raters to properly evaluate the results without time pressure. A binomial statistical test of the user study results, as presented in Table 4, suggests that our results are statistically significant (p-value $<5\%$).
Appendix B Additional Experiments
In Section B.1, we start by providing additional comparisons and results of our method. Then, in Section B.2, we present experiments on using different perceptual metrics. Following that, in Section B.4, we test the effect of different sizes for vital layer set. Next, in Section B.5, we provide latent nudging experiments. Furthermore, in Section B.6 we present a full visualization of our layer bypassing method. Finally, in Section B.7, we test our method on the Stable Diffusion 3 backbone.
B.1 Additional Comparisons and Results
Input SDEdit [53] P2P+NTI [35, 54] Instruct-P2P [18] MagicBrush [97] MasaCTRL [19] Stable Flow (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a bear with a long hair”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a man raising his hand”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a sheep with a yellow hat”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a cat wearing purple sunglasses”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a chicken during sunset”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of an elephant next to a blue ball”
Figure 14: Baselines Qualitative Comparison on Automatic Dataset. As explained in Section 4.1, we compare our method against the baselines on real images extracted from the COCO [49] dataset. We find that SDEdit [53] struggles with preserving the object identities and backgrounds (e.g., bear and chicken examples). P2P+NTI  [35, 54] struggles with preserving object identities (e.g., bear and person examples) and with adding new objects (e.g., missing hat in the sheep example and missing ball in the elephant example). Instruct-P2P [18] and MagicBrush [97] struggle with non-rigid editing (e.g., person raising hand). MasaCTRL [19] struggles with preserving object identities (e.g., bear and person examples) and adding new objects (e.g., sheep and cat examples). Our method, on the other hand, is able to adhere to the editing prompt while preserving the identities.
Input (1) Inj. all layers (2) Inj. non-vital layers (3) Extension all layers (4) w/o latent nudging Stable Flow (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a bear with a long hair”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a man raising his hand”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a sheep with a yellow hat”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a cat wearing purple sunglasses”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of a chicken during sunset”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photo of an elephant next to a blue ball”
Figure 15: Ablations Qualitative Comparison on Automatic Dataset. As explained in Section 4.3, we compare our method against several ablation cases on real images extracted from the COCO [49] dataset. As can be seen, we found that (1) performing attention injection in all the layers or performing (3) an attention extension in all the layers encourages the model to directly copy the input image while neglecting the target prompt. In addition, (2) performing an attention extension in the non-vital layers or (4) removing the latent nudging reduces the input image similarity significantly.
Refer to caption Refer to caption Refer to caption Refer to caption
Input “A ‘Stable Flow’ neon sign” “A ‘P = NP’ neon sign” “A neon sign of avocados”
Refer to caption Refer to caption Refer to caption Refer to caption
Input “A wooden lion” “A wooden toilet” “A wooden noodles bowl”
Refer to caption Refer to caption Refer to caption Refer to caption
Input “A hedgehog” “A shark” “A bird”
Refer to caption Refer to caption Refer to caption Refer to caption
Input “Jumping” “Sitting” “Putting its paw on a stone”
Figure 16: Additional Results. We provide various editing results of our method. These different edits are done using the same vital layer set.
Refer to caption Refer to caption Refer to caption Refer to caption
Input “An albino porcupine” “A horse” “A crow”
Refer to caption Refer to caption Refer to caption Refer to caption
Input “Wearing a red shirt” “Wearing purple jeans” “Wearing glasses”
Refer to caption Refer to caption Refer to caption Refer to caption
Input “The text ‘FLUX’ is written “A camel in the background” “A cat inside the bag”
on the bag”
Refer to caption Refer to caption Refer to caption Refer to caption
Input “A pink car” “A man driving the car” “In the evening”
Figure 17: Additional Results. We provide various editing results of our method. These different edits are done using the same vital layer set.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Input “Minds” “Think” “Alike”
Figure 18: Additional Results. Given an input image that contain a text, our method cat edit the text while keeping the background and style.
In Figure 14 we provide an additional qualitative comparison of our method against the baselines on real images extracted from the COCO [49] dataset, as explained in Section 4.1. As can be seen, SDEdit [53] struggles with preserving the object identities and backgrounds (e.g., the bear and chicken examples). P2P+NTI  [35, 54] struggles with preserving object identities (e.g., the bear and person examples) and with adding new objects (e.g., the missing hat in the sheep example and missing ball in the elephant example). Instruct-P2P [18] and MagicBrush [97] struggle with non-rigid editing (e.g., the person raising hand example). MasaCTRL [19] struggles with preserving object identities (e.g., the bear and person examples) and adding new objects (e.g., the sheep and cat examples). Our method, on the other hand, is able to adhere to the editing prompt while preserving the identities.
Next, in Figure 15, we provide a qualitative comparison of the ablated cases that are explained in, Section 4.3. As can be seen, we found that (1) performing attention injection in all the layers or performing (3) an attention extension in all the layers, encourages the model to directly copy the input image while neglecting the target prompt. In addition, (2) performing an attention extension in the non-vital layers or (4) removing the latent nudging reduces the input image similarity significantly.
Finally, in Figures 16 and 17, we present additional image editing results using our method.
B.2 Different Perceptual Metrics
00555510101010151515152020202025252525303030303535353540404040454545455050505055555555606060600.920.920.920.920.940.940.940.940.960.960.960.960.980.980.980.98Layer IndexCLIP Perceptual Similarity
Figure 19: Layer Removal Quantitative Comparison Using CLIP. As explained in Section B.2, we measured the effect of removing each layer of the model by calculating the CLIP [68] perceptual similarity between the generated images with and without this layer. Lower perceptual similarity indicates significant changes in the generated images. As can be seen, removing certain layers significantly affects the generated images, while others have minimal impact. Importantly, influential layers are distributed across the transformer rather than concentrated in specific regions. Note that the first vital layers were omitted for clarity (as their perceptual similarity approached zero).
00555510101010151515152020202025252525303030303535353540404040454545455050505055555555606060600.80.80.80.80.850.850.850.850.90.90.90.90.950.950.950.95Layer IndexDINOv1 Perceptual Similarity
Figure 20: Layer Removal Quantitative Comparison Using DINOv1. As explained in Section B.2, we measured the effect of removing each layer of the model by calculating the DINOv1 [20] perceptual similarity between the generated images with and without this layer. Lower perceptual similarity indicates significant changes in the generated images. As can be seen, removing certain layers significantly affects the generated images, while others have minimal impact. Importantly, influential layers are distributed across the transformer rather than concentrated in specific regions. Note that the first vital layers were omitted for clarity (as their perceptual similarity approached zero).
0055551010101015151515202020202525252530303030353535354040404045454545505050505555555560606060111111111111111111111111Layer Index(1 - LPIPS) Perceptual Similarity
Figure 21: Layer Removal Quantitative Comparison Using LPIPS. As explained in Section B.2, we measured the effect of removing each layer of the model by calculating the (1 - LPIPS) [98] perceptual similarity between the generated images with and without this layer. Lower perceptual similarity indicates significant changes in the generated images. As can be seen, removing certain layers significantly affects the generated images, while others have minimal impact. Importantly, influential layers are distributed across the transformer rather than concentrated in specific regions. Note that the first vital layers were omitted for clarity (as their perceptual similarity approached zero).
Input DINO [20, 58] CLIP [68] LPIPS [98]
Refer to caption Refer to caption Refer to caption Refer to caption
“A man holding a cup of tea”
Refer to caption Refer to caption Refer to caption Refer to caption
“An avocado at the beach”
Refer to caption Refer to caption Refer to caption Refer to caption
“A blackboard with the text “Vital Layers Are All You Need” ”
Refer to caption Refer to caption Refer to caption Refer to caption
“The floor is covered with snow”
Figure 22: Metrics Qualitative Comparison. As described in Section B.2, we also experimented with other perceptual metrics. We found DINOv2 [58] and DINOv1 [20] to produce the same set of vital layer. While CLIP [68] and LPIPS [98] replaced two layers in the vital layers set (though they include most of the vital layer set as in DINO). As can be seen, the differences between these sets are negligible when editing images.
As explained in Section 3.1, we assess the impact of each layer by measuring the perceptual similarity between $G_{\textit{ref}}$ and $G_{\ell}$ using DINOv2 [58]. It raises the question of the importance of the specific perceptual [43] similarity metric when determining the vital layers.
To this end, we also experiment with different perceptual metrics: DINOv1 [20], CLIP [68], and LPIPS [98]. In Figures 19, 20 and 21 we plot the perceptual similarity per layer for each of these metrics. The vital layers, ordered by vitality, as defined in Equation 1, for each metric are:
  • DINOv2 — $[1,0,2,18,53,28,54,17,56,25]$.
  • DINOv1 — $[1,0,2,18,53,56,54,25,28,17]$.
  • CLIP — $[2,0,1,18,53,56,54,4,17,3]$.
  • LPIPS — $[0,1,2,18,17,56,53,54,6,4]$.
As can be seen, the vital set $V$ is equivalent for DINOv2 and DINOv1 (even though there is a disagreement on the order). In addition, all the metrics include the set of $\{1,0,2,18,53,54,17,56\}$ to be included in the vital set, while DINOv1 and DINOv2 suggest also including $\{28,25\}$, CLIP suggests including $\{3,4\}$ instead and LPIPS suggests including $\{6,4\}$ instead. In Figure 22 we edited images with these slightly different vital layer sets, and found the differences to be negligible in practice.
B.3 VLM-Based Quantiative Metric
Table 5: VLM-Based quantitative comparison. For each method, we used Phi-3.5-vision [3] VLM to compute the percentage of editing results that follow the text prompt and of the results that change only the essential parts of the image. P2P+NTI [35, 54], Instruct-P2P [18], and MasaCTRL [19] suffer from low similarity to the text prompt. SDEdit [94] and MagicBrush [97] adhere more to the text prompt, but they struggle with avoiding unintended changes.
Method Text Following ()($\uparrow$)( ↑ ) Modify only essential ()($\uparrow$)( ↑ ) SDEdit [88] 86.66% 21.66% P2P+NTI [33, 51] 68.33% 26.66% Instruct-P2P [17] 33.33% 26.66% MagicBrush [91] 88.33% 46.66% MasaCTRL [18] 33.33% 06.66% Stable Flow (ours) 83.33% 61.66%
As explained in Section 4.1, We evaluated the editing results using three widely-used CLIP-based metrics: $\text{CLIP}_{img}$, $\text{CLIP}_{txt}$, and $\text{CLIP}_{dir}$. In addition, we experimented with a VLM-based metric using the Phi-3.5-vision [3] VLM that was trained specifically for the task of multiple image comparison. For each input image $x$, editing prompt $p$, and editing result $\hat{x}$, we computed the following two metrics: (1) Text Following — we presented the VLM the edited image $\hat{x}$ and the editing prompt $p$ and asked it “Does this image correspond to the text $p$? Answer yes or no.”. (2) Modify only essential parts — we extracted the prompt instruction and presented it, along with $x$ and $\hat{x}$, to the VLM and asked it “Is the only difference between these two images the text PROMPT? Answer yes or no”. For each metric, we calculated the number of times that the VLM answered “yes”. As demonstrated in Table 5, the VLM-based metric follows the same trend as the CLIP-based metrics:P2P+NTI [35, 54], Instruct-P2P [18], and MasaCTRL [19] suffer from low similarity to the text prompt. SDEdit [94] and MagicBrush [97] adhere more to the text prompt, but they struggle with avoiding unintended changes.
B.4 Number of Vital Layers
Input $100\%$ $80\%$ $60\%$ $40\%$ $20\%$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A man holding a cup of tea”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“An avocado at the beach”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A blackboard with the text “Vital Layers Are All You Need” ”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“The floor is covered with snow”
Figure 23: Number of Vital Layers Comparison. As explained in Section B.4, we experimented with choosing a different portion of the calculated vital layer set $V$. As can be seen, when removing $20\%$ of the vital layer set, the changes are negligible. However, when removing more than that, the editing results include unintended changes, such as identity changes (e.g., the man and woman examples) and background changes (e.g., the cat and blackboard examples).
The somewhat agnostic nature of our method to the specific perceptual metric, as described in Section B.2, raises the question of the importance of the entire vital layer set $V$ to the editing task. To this end, in Figure 23 we experimented with omitting a growing number of vital layers and testing the editing results. As can be seen, when removing $20\%$ of the vital layer set, the changes are negligible. However, when removing more than that, the editing results include unintended changes, such as identity changes (e.g., man and woman examples) and background changes (e.g., cat and blackboard examples). This is consistent with the results from Section B.2 that show that the least vital layers for each perceptual metric are less important for the image editing task.
B.5 Latent Nudging Experiments
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input $\lambda=1.0$ $\lambda=1.1$ $\lambda=1.15$ $\lambda=3.0$
Refer to caption Refer to caption Refer to caption Refer to caption
$\lambda=0.9$ $\lambda=0.8$ $\lambda=0.7$ $\lambda=0.6$
Figure 24: Latent Nudging Values. As described in Section B.5, we empirically tested different values for the latent nudging hyperparameter $\lambda$. In our experiments, we performed inversion using the inverse Euler ODE solver with a high number of 1,000 inversion (and denoising) steps, to reduce the inversion error. However, even when using such a high number of inversion/denoising steps, we notice that when not using latent nudging (i.e., $\lambda=1.0$), the reconstruction quality is poor (notice the eyes and the legs of the dog). Next, we found that $\lambda=1.15$ is the smallest value that enables full reconstruction using the inverse Euler solver. Furthermore, nudging values that are too high (e.g., $\lambda=3.0$) result in saturated images. Lastly, we notice that reducing nudging values ($\lambda<1.0$) severely damages the reconstruction quality.
Input Reconstruction Caching only Latent Nudging
Refer to caption Refer to caption Refer to caption Refer to caption
“A rabbit toy sitting and wearing
pink socks during the late afternoon”
Refer to caption Refer to caption Refer to caption Refer to caption
“The cat is yelling and raising its paw”
Refer to caption Refer to caption Refer to caption Refer to caption
“A rubber duck next to a purple ball during a sunny day”
Refer to caption Refer to caption Refer to caption Refer to caption
“A man with a long hair”
Figure 25: Latent Caching. As explained in Section B.5, we also tested a latent caching approach [12], in which we saved the series of latents during the inversion process without applying latent nudging. As can be seen, this approach indeed achieves perfect inversion (second column), but (third column) still struggles with preserving the identities while editing the image (e.g., the rabbit and duck examples) or significantly alters the image (e.g., the cat and man examples). On the other hand, our method with the latent nudging (fourth column) is able to preserve the identities during editing.
As described in Section 3.3 of the main paper, we proposed using a latent nudging technique to avoid the bad reconstruction quality of vanilla inverse Euler ODE solver. We suggest multiplying the initial latent $z_{0}$ by a small scalar $\lambda=1.15$ to slightly offset it from the training distribution. As shown in Figure 24, we empirically tested different values for the latent nudging hyperparameter $\lambda$. We performed inversion using the inverse Euler ODE solver with a high number of 1,000 inversion (and denoising) steps, to reduce the inversion error. However, even when using such a high number of inversion/denoising steps, we notice that when not using latent nudging (i.e., $\lambda=1.0$), the reconstruction quality is poor (notice the eyes and the legs of the dog). Next, we found that $\lambda=1.15$ is the smallest value that enables full reconstruction using the inverse Euler solver. Furthermore, nudging values that are too high (e.g., $\lambda=3.0$) result in saturated images. Lastly, we notice that decreasing nudging values (i.e., $\lambda<1.0$) severely damages the reconstruction quality.
In addition, we experiment with a simpler inversion variant based on latent caching (termed DDPM bucketing in DiffUHaul [12]), in which we saved the series of latents during the inversion process without applying latent nudging. As shown in Figure 25, this approach indeed achieves perfect inversion (second column), but (third column) still struggles with preserving the identities while editing the image (e.g., the rabbit and duck examples) or significantly alters the image (e.g., the cat and man examples). On the other hand, our method (fourth column) with the latent nudging is able to preserve the identities during editing. In practice, we found that using latent caching in addition to latent nudging enables inversion with a lower number of steps (50 steps), hence, this is the approach we used.
B.6 Layer Bypassing Visualization
Refer to caption Refer to caption Refer to caption
$G_{\textit{ref}}$ $G_{0}$ $G_{1}$
Refer to caption Refer to caption Refer to caption
$G_{2}$ $G_{3}$ $G_{4}$
Refer to caption Refer to caption Refer to caption
$G_{5}$ $G_{6}$ $G_{7}$
[…]
Refer to caption Refer to caption Refer to caption
$G_{\textit{ref}}$ $G_{8}$ $G_{9}$
Refer to caption Refer to caption Refer to caption
$G_{10}$ $G_{11}$ $G_{12}$
Refer to caption Refer to caption Refer to caption
$G_{13}$ $G_{14}$ $G_{15}$
[…]
Refer to caption Refer to caption Refer to caption
$G_{\textit{ref}}$ $G_{16}$ $G_{17}$
Refer to caption Refer to caption Refer to caption
$G_{18}$ $G_{19}$ $G_{20}$
Refer to caption Refer to caption Refer to caption
$G_{21}$ $G_{22}$ $G_{23}$
[…]
Refer to caption Refer to caption Refer to caption
$G_{\textit{ref}}$ $G_{24}$ $G_{25}$
Refer to caption Refer to caption Refer to caption
$G_{26}$ $G_{27}$ $G_{28}$
Refer to caption Refer to caption Refer to caption
$G_{29}$ $G_{30}$ $G_{31}$
[…]
Refer to caption Refer to caption Refer to caption
$G_{\textit{ref}}$ $G_{32}$ $G_{33}$
Refer to caption Refer to caption Refer to caption
$G_{34}$ $G_{35}$ $G_{36}$
Refer to caption Refer to caption Refer to caption
$G_{37}$ $G_{38}$ $G_{39}$
[…]
Refer to caption Refer to caption Refer to caption
$G_{\textit{ref}}$ $G_{40}$ $G_{41}$
Refer to caption Refer to caption Refer to caption
$G_{42}$ $G_{43}$ $G_{44}$
Refer to caption Refer to caption Refer to caption
$G_{45}$ $G_{46}$ $G_{47}$
[…]
Refer to caption Refer to caption Refer to caption
$G_{\textit{ref}}$ $G_{48}$ $G_{49}$
Refer to caption Refer to caption Refer to caption
$G_{50}$ $G_{51}$ $G_{52}$
Refer to caption Refer to caption Refer to caption
$G_{53}$ $G_{54}$ $G_{55}$
[…]
Refer to caption Refer to caption
$G_{\textit{ref}}$ $G_{56}$
Figure 33: Full Layer Bypassing Visualization for Flux. We visualize the individual layer bypassing study we conducted, as described in Section B.6. We start by generating a set of images $G_{\textit{ref}}$ using a fixed set of seeds and prompts. Then, we bypass each layer $\ell$ by using its residual connection and generate the set of images $G_{\ell}$ using the same fixed set of prompts and seeds. In this visualization, $G_{56}$ is a vital layer.
As explained in Section 3.1, to quantify layer importance in FLUX model, we devised a systematic evaluation approach. Using ChatGPT [57], we automatically generated a set $P$ of $k=64$ diverse text prompts, and draw a set $S$ of random seeds. Each of these prompts was used to generate a reference image, yielding a set $G_{\textit{ref}}$. For each DiT layer $\ell\in\mathbb{L}$, we performed an ablation by bypassing the layer using its residual connection. This process generated a set of images $G_{\ell}$ from the same prompts and seeds.
[…]
B.7 Stable Diffusion 3 Results
005555101010101515151520202020000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8Layer IndexPerceptual Similarity
Figure 34: Layer Removal Quantitative Comparison Stable Diffusion 3. As explained in Section B.7, we measured the effect of removing each layer of the model by calculating the perceptual similarity between the generated images with and without this layer. Lower perceptual similarity indicates significant changes in the generated images. As can be seen, removing certain layers significantly affects the generated images, while others have minimal impact. For a visual comparison, please refer to Figure 35.
$G_{\textit{ref}}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{0}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{1}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{7}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{8}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{9}$
Refer to caption Refer to caption Refer to caption Refer to caption
$G_{21}$
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 35: Layer Removal Qualitative Comparison Stable Diffusion 3. As explained in Section B.7, we illustrate the qualitative differences between vital and non-vital layers. While bypassing non-vital layers ($G_{1}$ and $G_{21}$) results in modest alterations, bypassing vital layers leads to significant changes: complete noise generation ($G_{0}$), or severe distortions ($G_{7}$, $G_{8}$ and $G_{9}$). For a quantitative comparison, please refer to Figure 34
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input “A dog statue” “A lemur” “A pig” “A gecko”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input “Angry” “Closing his eyes” “Wearing glasses” “An old man”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input “A pink car” “A blue car” “An orange car” “A green car”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input “Closing its eyes” “Wearing a red hat” “Wearing green “Next to a purple stone”
glasses”
Figure 36: Stable Diffusion 3 Editing Results. As explained in Section B.7, we tested our Stable Flow method on the Stable Diffusion 3 backbone [26]. As can be seen, we are able to perform various editing operations using the same mechanism of injecting the reference image information into the vital layers of the model.
All the experiments in the main paper were based on the FLUX.1-dev [47] model. We also experimented with a different DiT text-to-image flow model named Stable Diffusion 3 [26] based on the Diffusers [86] implementation of the medium model.
As described in Section 3.1, we measured the importance of each of the layers of this model. As shown in Figure 34, we measured the effect of removing each layer from the model by calculating the perceptual similarity between the generated images with and without this layer. Lower perceptual similarity indicates significant changes in the generated images. As can be seen, removing certain layers significantly affects the generated images, while others have minimal impact.
Next, in Figure 35 we illustrate the qualitative differences between vital and non-vital layers. While bypassing non-vital layers ($G_{1}$ and $G_{21}$) results in modest alterations, bypassing vital layers leads to significant changes: complete noise generation ($G_{0}$) or severe distortions ($G_{7}$, $G_{8}$, and $G_{9}$).
Finally, in Figure 36, we perform various editing operations using the same mechanism of injecting the reference image information into the vital layers of the model, as described in Section 3.2.